### v1

@@ 4.5 Simplification with Zero Rest Twist
While capturing the twisting effects of elastic strands is essential for visual fidelity, it is relatively rare to exactly specify initial twist or edge angles when designing a strand geometry, e.g., for hair assets, because edge angles have almost no effect on the visualization result at the static state, and capturing and specifying edge angle information is challenging [Luo2013hair, Hu2015hair]. In addition, some elastic rod simulators do not necessarily support edge angles [Selle2008mass, Umetani2014rod, Shi2023curls] and it is generally difficult for humans to intuit how the edge angles influence the strand dynamics for simulators that do consider the edge angles [Bergou2008, Bergou:2010:DVT:1778765.1778853, Kugelstadt2016rod, Soler2018rod]. As a result, it is common to specify only vertex positions and initialize edge angles $\bm{\theta}$ to zero. Given such zero edge angles, which also yield zero initial twist $\mathbf{m}$, setting the rest twist $\mathbf{\bar{m}}$ to zero completely eliminates the twist energy (16) and its gradient (17). As such, we can potentially exclude $\mathbf{\bar{m}}$ from $\mathbf{\bar{s}}$, leading to an effective DOF count of $(5N-10)$ for $\mathbf{\bar{s}}$ and thus reducing the computational cost for rest shape optimization.
In practice, rest shape optimization without the rest twist still works well (i.e., achieves a stable static equilibrium) if the bending of the strands introduces only a limited amount of twisting given the tight coupling of bending and twisting in the DER formulation. Such cases do occur, so we will evaluate and compare the rest shape optimization with and without rest twist in terms of performance and quality.
@@ 5 Results and Discussions
We implemented our method in C++20 with double-precision floating-point for scalar values. When multiple strands are involved, we parallelize forward simulation and inverse problems with OpenMP, processing each strand in parallel. We executed the examples on a desktop machine with an Intel Core i7-9700 (8 cores) with 16GB RAM. We use 60 frames per second with a single simulation step except for Figure  Author Guidelines for EUROGRAPHICS Proceedings Manuscripts and Figure 9, where we used five steps. We include the rest twist in the optimization unless otherwise mentioned. We set $\rho=10^{3}\ \mathrm{kg/m^{3}}$, $r=10^{-3}\ \mathrm{m}$, $c_{\mathrm{stretch}}=10^{8}\ \mathrm{kg/(m\cdot s^{2})}$, $c_{\mathrm{bend}}=10^{8}\ \mathrm{kg/(m\cdot s^{2})}$, and $c_{\mathrm{twist}}=10^{8}\ \mathrm{kg/(m\cdot s^{2})}$ unless otherwise specified. For visualization purposes, we enlarge the vertices and edges. The first two black vertices and cyan edge indicate the “fixed” vertices and edge, respectively, while the red vertices and blue and white edges represent movable vertices and edges with active DOFs, respectively (except for Figure  Author Guidelines for EUROGRAPHICS Proceedings Manuscripts and Figure 9). In addition, we render the material frames at the centers of the corresponding edges as green and magenta lines whose lengths are the same as the rest lengths of the edges.
@@ 5.1 Single Vertical Strand
To illustrate the effectiveness of our rest shape optimization, we first experiment with a single vertical elastic strand, uniformly discretized with 20 vertices, as shown in Figure 2. With three different stretch coefficients $c_{\mathrm{stretch}}$ ($5\times 10^{3},5\times 10^{4}$, and $5\times 10^{5}$), we compare simulation results with the following three cases:
@@
Figure 3: Convergence profile for GD, Newton, and GN (ours). Both Newton and GN quickly converge while GD requires much more iterations.
@@ 5.2 Nonlinear Optimizer Comparisons
For Newton’s method, we compute the exact Hessian for (37) with the stretch force (since bending and twisting will not be introduced in this setting) and ensure an SPD system via Hessian projection per edge. Figure 3 shows a convergence profile of the objective over optimization iterations. The back-tracking line search leads GD to require tiny steps due to the stiff system, and GD fails to sufficiently decrease the objective within 500 iterations. Given the sparse structure of the system, linear solves (required only for Newton-type optimizers, but not for GD) with a sparse Cholesky solver are sufficiently fast, and thus GN is much faster (GD took $10.7$ s with 500 iterations whereas GN took $0.11$ s with 6 iterations). With Newton’s method, we do not observe a particular benefit on convergence over Gauss-Newton while computing the exact Hessian introduces an additional cost and significant implementation complexity associated with the third order tensors. These observations make GN most suitable as our optimizer.
(a) $c_{\mathrm{bend}}=10^{8}$
(b) $c_{\mathrm{bend}}=10^{9}$
(c) $c_{\mathrm{bend}}=10^{10}$
Figure 4: Evaluation with a single horizontal strand. For each $c_{\mathrm{bend}}$ trio: (left) simulation results with naive initialization, (middle) with rest shape optimization, and (right) with rest shape optimization and zero gravity during simulation. Our optimizer modifies the rest curvature more significantly for softer strands to achieve static equilibrium.
@@ 5.3 Single Horizontal Strand
Next, we test with a horizontal strand uniformly discretized with 20 vertices, as shown in Figure 4. With three different bending coefficients $c_{\mathrm{bend}}$ ($10^{8},10^{9}$, and $10^{10}$), we compare the three cases:
@@
Due to the horizontal arrangement of the strands, our rest shape optimizer modifies the rest curvatures only (but not rest lengths or rest twists). We can achieve static equilibrium with all three different coefficients, with our optimizer modifying the rest curvature more significantly for the softer materials. While increasing the stiffness is one way to reduce the sagging at equilibrium, this approach can change the strand dynamics and our perception of the material.
@@ 5.4 External Load
We also introduce an additional external load (which is a force 10×10$\times$10 × larger than gravity) exactly on the non-root-end vertex of the horizontal strand, as shown in Figure 5. While static equilibrium can be achieved even with the load, our method modifies the rest curvature further to cancel the force due to the load.
@@ 5.5 Stress Test
To demonstrate the capabilities of our method with high-resolution strands, we experiment with a horizontal strand discretized with $1,000$ vertices, as shown in Figure 6 (notably, hair strands are more commonly discretized with up to $100$ vertices [Shi2023curls]). Given the horizontal structure of the strand, bending does not introduce twisting, and thus we can safely exclude the rest twist from the rest shape optimization. Our method achieved static equilibrium in both cases, i.e., including and excluding rest twist for the rest shape optimization. The computation times were $0.087$ s and $0.072$ s which are proportional to the number of DOFs, $(6N-12)$ and $(5N-10)$, respectively.
@@ 5.6 Single Hair-Like Strand
Next, we experiment with a single hair-like strand (which can introduce non-negligible torsion effects, unlike vertical and horizontal strands), as shown in Figure 7, to evaluate our method including/excluding rest twist in the optimization, along with the global-local initialization [Hsu2022sag]. Specifically, we compare the following schemes:
@@
    Global-local initialization: [Hsu2022sag].
Since directly applying the global-local initialization [Hsu2022sag] to DER is not possible, as discussed in Sec. 2.2.3, we use the following two simplifications. First, we use the minimal stencil of the bending and twisting as a unified single force element to form a global linear system, assuming that the force element is linear/angular momentum preserving. Second, we first determine rest length from the stretch force without considering the bending and twisting forces to avoid transforming the local step into a global problem assuming relatively small changes in the rest length due to the large $c_{\mathrm{stretch}}$. Then, we optimize the rest curvature and rest twist using the rest length and local-element forces (computed by the global linear solve) in a least-squares way.
The strand with non-zero rest twist achieves static equilibrium and retains the same shape. By contrast, while the strand with zero rest twist can achieve static equilibrium at first, it is relatively unstable and flips in its middle part to reach a more stable static equilibrium. This is because the bending forces need to cancel any torques by themselves (as the twist forces are zero by construction at first) leading to significant changes in the rest curvatures and thus the unstable static equilibrium. Due to the necessary simplifications, the global-local initialization [Hsu2022sag] failed to achieve static equilibrium (although the global-local initialization, as described above, was able to achieve static equilibrium for perfectly vertical or horizontal strands in our experiments).
@@ 5.7 Box Constraint Evaluation
Without using the box constraints, the rest shape parameters can be modified significantly, leading to a relatively unstable static equilibrium and thus failing to retain the original strand shape. In particular, for softer materials, the rest shape change can be excessive, introducing stability issues in the forward simulation. By contrast, the use of box constraints can avoid introducing stability problems and achieve stable strand dynamics and states, preserving the original shape better compared to the naive initialization.
@@
Figure 9: Hair simulation with complex strand geometry. Hair strands sag with the naive initialization while our method enables hair strands to retain their original shapes, exhibit natural motions due to the prescribed move of the root vertices, and then return toward the original shape.
@@ 5.8 Evaluation with Complex Strand Geometry
To evaluate the effectiveness of our method in more complex scenarios, we experiment with hair strand data released publicly by Hu et al. [Hu2015hair]. We use $c_{\mathrm{stretch}}=3.0\times 10^{8},c_{\mathrm{bend}}=3.0\times 10^{8}$, and $c_{\mathrm{twist}}=3.0\times 10^{8}$. Figure 9 compares our method with the naive initialization. While the hair strands sag due to the gravity with the naive initialization, the strands with our method retain the original hair style. When the root vertices are rotated in a prescribed way, both approaches generate natural and comparable hair motions. After the root vertices are stopped at the same position as at the start, hair strands with our approach retain the original hair shape, unlike those with the naive initialization. Figure  Author Guidelines for EUROGRAPHICS Proceedings Manuscripts shows hair simulations using our method with another complex hair style.

### v2

@@ 4.5 Simplification with Zero Rest Twist
While capturing the twisting effects of elastic strands is essential for visual fidelity, it is relatively rare to exactly specify initial twist or edge angles when designing a strand geometry, e.g., for hair assets, because edge angles have almost no effect on the visualization result at the static state, and capturing and specifying edge angle information is challenging [LLR13, HMLL15]. In addition, some elastic rod simulators do not necessarily support edge angles [SLF08, USS14, SWP23] and it is generally difficult for humans to intuit how the edge angles influence the strand dynamics for simulators that do consider the edge angles [BWR08, BAV10, KS16, SMSH18]. As a result, it is common to specify only vertex positions and initialize edge angles $\bm{\theta}$ to zero. Given such zero edge angles, which also yield zero initial twist $\mathbf{m}$, setting the rest twist $\mathbf{\bar{m}}$ to zero completely eliminates the twist energy (16) and its gradient (17). As such, we can potentially exclude $\mathbf{\bar{m}}$ from $\mathbf{\bar{s}}$, leading to an effective DOF count of $(5N-10)$ for $\mathbf{\bar{s}}$ and thus reducing the computational cost for rest shape optimization.
In practice, rest shape optimization without the rest twist still works well (i.e., achieves a stable static equilibrium) if the bending of the strands introduces only a limited amount of twisting given the tight coupling of bending and twisting in the DER formulation. Such cases do occur, so we will evaluate and compare the rest shape optimization with and without rest twist in terms of performance and quality (see Sec. 5.5 and Sec. 5.6).
4.6 Discussion of Norm Formulations
Eliding unnecessary coefficients for a simpler analysis, the objective based on the $L^{2}$ norm of the generalized force $\mathbf{f}$ is given by $\left\lVert\mathbf{f}\right\rVert_{2}^{2}$ [TKA11] whereas our kinetic-energy-based objective is $\left\lVert\mathbf{f}\right\rVert_{\mathbf{M}^{-1}}^{2}$ due to (20). As an example, consider the inertia force (23) due to generalized gravity $\mathbf{g}$ since the inertia is the unique trigger for non-zero elastic forces (which are expected to cancel the inertia) in our framework. In this case, $\mathbf{f}=\mathbf{M}\mathbf{g}$, and thus we get $\mathbf{g}^{T}\mathbf{M}^{2}\mathbf{g}$ for the force-norm-based objective versus $\mathbf{g}^{T}\mathbf{M}\mathbf{g}$ for ours. As such, when mass/inertia of the strands is too small or large, squaring $\mathbf{M}$ can negatively impact the numerical conditioning of the system. In practice, Gauss-Newton optimization on $\left\lVert\mathbf{f}\right\rVert_{2}^{2}$ through $(\mathbf{J}^{T}\mathbf{J})\Delta\mathbf{\bar{s}}=-\mathbf{J}^{T}\mathbf{f}$ (where $\mathbf{J}$ denotes the Jacobian of $\mathbf{f}$ with respect to $\mathbf{\bar{s}}$) fails to reach sufficiently small $\left\lVert\mathbf{f}\right\rVert_{2}^{2}$ because $\mathbf{J}^{T}\mathbf{J}$ has entries significantly different in their magnitude and thus is numerically ill-conditioned, neglecting some force components. Consequently, extended precision would be required for the optimization to consistently succeed [TKA11]. Additionally, it is worth noting that our formulation properly accounts for the energy contributions from each component (unlike simple scaling and dimension analysis approaches that merely mitigate numerical precision problems) and is essential for DER as it involves variables in different dimensions (vertex positions and edge angles).
@@ 5 Results and Discussions
We implemented our method in C++20 with double-precision floating-point for scalars. When multiple strands are involved, we parallelize forward simulation and inverse problems with OpenMP, processing each strand in parallel. We executed the examples on a desktop machine with an Intel Core i7-9700 (8 cores) with 16GB RAM. We use 60 frames per second with a single simulation step except for Figure Rest Shape Optimization for Sag-Free Discrete Elastic Rods and Figure 11, where we used five steps. We include the rest twist in the optimization unless otherwise mentioned. We set $\rho=10^{3}\ \mathrm{kg/m^{3}}$, $r=10^{-3}\ \mathrm{m}$, $c_{\mathrm{stretch}}=10^{8}\ \mathrm{kg/(m\cdot s^{2})}$, $c_{\mathrm{bend}}=10^{8}\ \mathrm{kg/(m\cdot s^{2})}$, and $c_{\mathrm{twist}}=10^{8}\ \mathrm{kg/(m\cdot s^{2})}$ unless otherwise specified. For visualization purposes, we enlarge the vertices and edges. The first two black vertices and cyan edge indicate the “fixed” vertices and edge, respectively, while the red vertices and blue and white edges represent movable vertices and edges with active DOFs, respectively (except for Figures Rest Shape Optimization for Sag-Free Discrete Elastic Rods, 10, and 11). In addition, we render the material frames at the centers of the corresponding edges as green and magenta lines whose lengths are the same as the rest lengths of the edges.
@@ 5.1 Single Vertical Strand
To illustrate the effectiveness of our rest shape optimization, we first experiment with a single vertical elastic strand (whose length is $1\mathrm{m}$), uniformly discretized with 20 vertices, as shown in Figure 2. With three different stretch coefficients $c_{\mathrm{stretch}}$ ($5\times 10^{3},5\times 10^{4}$, and $5\times 10^{5}$), we compare simulation results with the following three cases:
@@
Figure 3: Convergence profile for GD, Newton, and GN (ours). Both Newton and GN quickly converge while GD requires many more iterations.
@@ 5.2 Nonlinear Optimizer Comparisons
For Newton’s method, we compute the exact Hessian via the third-order tensors for (37) with the stretching force (since bending and twisting will not be introduced in this setting) and ensure an SPD system via Hessian projection per edge. Figure 3 shows a convergence profile of the objective over optimization iterations.
The back-tracking line search leads GD to require tiny steps due to the stiff system, and GD fails to sufficiently decrease the objective within 500 iterations. Due to the one-dimensional strand structure, a sparse Cholesky solver with matrix reordering can solve the system (linear solve is required only for Newton-type optimizers, but not for GD) with an almost linear cost and is sufficiently fast (per-iteration cost for GN is less than 2×2$\times$2 × of GD), and thus GN is much faster (GD took $10.7$ s with 500 iterations whereas GN took $0.11$ s with 6 iterations). Given GN’s low per-iteration cost, it is likely to outperform quasi-Newton methods. With Newton’s method, we do not observe a particular benefit on convergence over GN while computing the exact Hessian introduces an additional cost and significant implementation complexity associated with the third-order tensors. These observations make GN most suitable as our optimizer.
(a) $c_{\mathrm{bend}}=10^{7}$
(b) $c_{\mathrm{bend}}=10^{8}$
(c) $c_{\mathrm{bend}}=10^{9}$
(d) $c_{\mathrm{bend}}=10^{10}$
Figure 4: Evaluation with a single horizontal strand. For each $c_{\mathrm{bend}}$ trio: (left) simulation results with naive initialization, (middle) with rest shape optimization, and (right) with rest shape optimization and zero gravity during simulation. Our optimizer modifies the rest curvature more significantly for softer strands to achieve static equilibrium (b), (c), and (d) while it fails for too soft material (a).
@@ 5.3 Single Horizontal Strand
Next, we test with a horizontal strand uniformly discretized with 20 vertices, as shown in Figure 4, and used $\Delta\bm{\kappa}=10$ to allow larger curvature changes. With four different bending coefficients $c_{\mathrm{bend}}$ ($10^{7},10^{8},10^{9}$, and $10^{10}$), we compare the three cases:
@@
Due to the horizontal arrangement of the strands, our rest shape optimizer modifies the rest curvatures only (but not rest lengths or rest twists). While we can achieve static equilibrium with the three stiffer strands, with our optimizer modifying the rest curvature more significantly for the softer materials, the softest strand fails because it requires much larger rest curvature changes (which are not permitted by the box constraints). While increasing the stiffness is one way to reduce the sagging at equilibrium, this approach can change the strand dynamics and our perception of the material.
@@ 5.4 External Load
We also introduce an additional external load (which is a force 10×10$\times$10 × larger than gravity) exactly on the tail-end vertex of the horizontal strand, as shown in Figure 5. While static equilibrium can be achieved even with the load, our method modifies the rest curvature further to cancel the force due to the load.
@@ 5.5 Stress Test
To demonstrate the capabilities of our method with high-resolution strands, we experiment with a horizontal strand discretized with $1,000$ vertices, as shown in Figure 6 (notably, hair strands are more commonly discretized with up to $100$ vertices [SWP23]). Given the horizontal structure of the strand, bending does not introduce twisting, and thus we can safely exclude the rest twist from the rest shape optimization. Our method achieved static equilibrium in both cases, i.e., including and excluding rest twist for the rest shape optimization. The computation times were $0.087$ s and $0.072$ s which are proportional to the number of DOFs, $(6N-12)$ and $(5N-10)$, respectively.
@@ 5.6 Single Hair-Like Strand
Next, we experiment with a single hair-like strand (which can introduce non-negligible torsion effects, unlike vertical and horizontal strands), as shown in Figure 7, to evaluate our method including/excluding rest twist in the optimization, along with the global-local initialization [HTYW22]. Specifically, we compare the following schemes:
@@
    Global-local initialization: [HTYW22].
Since directly applying the global-local initialization [HTYW22] to DER is not possible, as discussed in Sec. 2.2.3, we use the following two simplifications. First, we use the minimal stencil of the bending and twisting as a unified single force element to form a global linear system, assuming that the force element is linear/angular momentum preserving. Second, we first determine rest length from the stretch force without considering the bending and twisting forces to avoid transforming the local step into a global problem assuming relatively small changes in the rest length due to the large $c_{\mathrm{stretch}}$. Then, we optimize the rest curvature and rest twist using the rest length and local-element forces (computed by the global linear solve) in a least-squares way.
The strand with non-zero rest twist achieves static equilibrium and retains the same shape. By contrast, while the strand with zero rest twist can achieve static equilibrium at first, it is relatively unstable and flips in its middle part to reach a more stable static equilibrium. This is because the bending forces need to cancel any torques by themselves (as the twist forces are zero by construction at first) leading to significant changes in the rest curvatures and thus the unstable static equilibrium. Due to the necessary simplifications, the global-local initialization [HTYW22] failed to achieve static equilibrium (although the global-local initialization, as described above, was able to achieve static equilibrium for perfectly vertical or horizontal strands in our experiments).
@@ 5.7 Box Constraint Evaluation
Without using the box constraints, the rest shape parameters can be modified significantly, leading to a relatively unstable static equilibrium and thus failing to retain the original strand shape. In particular, for softer materials, the rest shape change can be excessive, introducing stability issues in the forward simulation. By contrast, the use of box constraints can avoid introducing stability problems and achieve stable strand dynamics and states (although the strand needs to compromise static equilibrium), preserving the original shape better compared to the naive initialization.
(a) Force norm
(b) Equilibration
(c) Kinetic energy norm
Figure 9: Experiment with a hair-like strand. Optimization with the force norm and equilibration approach fails to achieve static equilibrium whereas our method succeeds.
Table 1: Results at solver convergence for the experiment in Figure 9. $\mathrm{min}(\mathbf{A})$ and $\mathrm{max}(\mathbf{A})$ denote the minimum and maximum magnitude of non-zeros in the system matrix $\mathbf{A}$, respectively, $\sigma$ their ratio ($\mathrm{max}(\mathbf{A})/\mathrm{min}(\mathbf{A})$), $\left\lVert\mathbf{f}\right\rVert_{2}$ and $\left\lVert\mathbf{f}\right\rVert_{\mathbf{M}^{-1}}$ the norm of corresponding objectives.
Schemes $\mathrm{min}(\mathbf{A})$ $\mathrm{max}(\mathbf{A})$ $\sigma$ $\left\lVert\mathbf{f}\right\rVert_{2}$ $\left\lVert\mathbf{f}\right\rVert_{\mathbf{M}^{-1}}$
Force norm $6.0\times 10^{-9}$ $8.4\times 10^{8}$ $1.4\times 10^{17}$ $3.1\times 10^{-4}$ $1.2\times 10^{1}$
Equilibration $1.9\times 10^{-14}$ $1.0\times 10^{0}$ $5.4\times 10^{13}$ $3.1\times 10^{-4}$ $1.2\times 10^{1}$
Kinetic energy norm $1.4\times 10^{-2}$ $2.7\times 10^{13}$ $1.9\times 10^{15}$ $4.3\times 10^{-8}$ $6.3\times 10^{-6}$
5.8 Evaluation of Norm Formulations
To demonstrate the effectiveness of our norm formulation, we experiment with the hair-like strand, as shown in Figure 9. We compare the following schemes:
  1. 1.
    Force norm: force-norm formulation $\left\lVert\mathbf{f}\right\rVert_{2}^{2}$ [TKA11] which forms the Hessian via $\mathbf{J}^{T}\mathbf{J}$;
  2. 2.
    Equilibration: force-norm formulation [TKA11] with symmetric equilibration [Bun71] to improve numerical conditioning of the system via $\mathbf{D}\mathbf{J}^{T}\mathbf{J}\mathbf{D}$ ($\mathbf{D}$: diagonal scaling matrix);
  3. 3.
    Kinetic energy norm: our kinetic-energy-based norm $\left\lVert\mathbf{f}\right\rVert_{\mathbf{M}^{-1}}^{2}$ which forms the Hessian via $\mathbf{J}^{T}\mathbf{M}^{-1}\mathbf{J}$.
We ignore the constant scaling of objectives for readability. Table 1 summarizes the result of the experiment. Optimization with the force norm fails to achieve static equilibrium because the system has entries significantly different in magnitude and thus neglects some force components. While equilibrating the matrix improved the numerical conditioning of the system with smaller variations in the matrix entries (i.e., smaller $\sigma$), it did not help to further reduce the objective values within the limited numerical precision (giving a result almost identical to the one with the force norm), because it neglects the energy contributions from each force component. By contrast, our kinetic-energy-based objective properly considers the energy contributions in the system and sufficiently reduces both the objectives $\left\lVert\mathbf{f}\right\rVert_{2}^{2}$ and $\left\lVert\mathbf{f}\right\rVert_{\mathbf{M}^{-1}}^{2}$, thus achieving static equilibrium.
(a) Naive initialization
(b) With rest shape optimization
Figure 10: Evaluation using a coiled hair-like strand, with its root vertex perturbed. The naive initialization fails to achieve static equilibrium and suffers from the stability issue, whereas the rest shape optimization enables sag-free and stable simulation.
5.9 Evaluation with Coiled Hair-Like Strand
To demonstrate that our method can handle curly strands, we experiment with a coiled hair-like strand, as shown in Figure 10, using $c_{\mathrm{stretch}}=c_{\mathrm{bend}}=c_{\mathrm{twist}}=10^{9}$. By simultaneously optimizing the rest length, rest curvature, and rest twist, our method achieves static equilibrium. Additionally, when the root vertices are perturbed, the optimized rest shape prevents too much strand deformation and thus stabilizes the simulation in this example, whereas the strand with the naive initialization exhibits unnatural oscillations.
@@ 5.9 Evaluation with Coiled Hair-Like Strand
Figure 11: Hair simulation with hair assets. Hair strands sag with the naive initialization while our method enables hair strands to retain their original shapes, exhibit natural motions due to the prescribed motion of the root vertices, and then return toward the original shape.
(a) Initial condition
(b) Naive initialization
(c) With optimization
Figure 12: Experiment with thinner strands. With naive initialization, strands sag significantly. With rest shape optimization, strands still sag (though sagging is mitigated) and exhibit unnatural lifting.
@@ 5.10 Evaluation with Complex Strand Geometry
To evaluate the effectiveness of our method in more complex scenarios, we experiment with hair strand data released publicly by Hu et al. [HMLL15]. We use $c_{\mathrm{stretch}}=3.0\times 10^{8},c_{\mathrm{bend}}=3.0\times 10^{8}$, and $c_{\mathrm{twist}}=3.0\times 10^{8}$. Figure 11 compares our method with the naive initialization. While the hair strands sag due to the gravity with the naive initialization, the strands with our method retain the original hair style. When the root vertices are rotated in a prescribed way, both approaches generate natural and comparable hair motions. After the root vertices are stopped at the same position as at the start, hair strands with our approach retain the original hair shape, unlike those with the naive initialization. Figure Rest Shape Optimization for Sag-Free Discrete Elastic Rods shows hair simulations using our method with another complex hair style.
In Figure 12, we use the same setting as Figure 11, but with thinner strands (strand radius $r=10^{-4}$). In this case, rest shape optimization fails to retain the original hair shape (while the sagging is mitigated compared to naive initialization) because the strands cannot exert sufficiently strong forces to support themselves. Additionally, significant rest shape changes can lead to unstable strand configurations, causing unnatural hair lifting.